# Scatterless interferences: Delay of laminar-to-turbulent flow transition by a lattice of subsurface phonons

### v1

@@ Abstract
Turbulent flows in boundary layers generate significantly higher surface friction drag compared to laminar flows. Avoidance or delay of laminar-to-turbulent flow transition is therefore critical for efficient operation of air vehicles and water vessels. Here we present a passive phonons-based solution that is based on engineered intervention with the unstable flow fluctuations that trigger the transition. The intervention is introduced by a tailored frequency- and wavevector-dependent interaction between the flow and the bounding surface. The surface is homogenous and smooth with all the engineered constituents buried in the subsurface. The subsurface configuration comprises a wall-parallel lattice of wall-normal phononic structural units. Each discrete phononic subsurface unit is designed to respond to the flow in a negative phase as to inhibit the flow instabilities at the point of contact. The collective effect of the lattice formation of these units, in turn, offers a form of interferences that causes the kinetic energy of the flow instabilities to decay downstream to the location of the intervention, thus delaying the transition. This concept presents a new paradigm for the design of aircrafts and ships beyond the prevailing paradigm of streamlined shaping of the surfaces.
@@
The interaction of a flowing fluid with a surface is a dynamical process that lies at the heart of a wide range of both natural and artificial systems. Two most consequential examples are the flow over an aircraft wing or around the hull of a ship. The nature of the fluid-structure interaction in these cases dictates the intensity of skin-friction drag, the location of onset of laminar-to-turbulent flow transition, and the possibility of flow separation--all with far-reaching implications on the overall performance of the vehicle [1, 2]. Skin friction drag in particular has profound economic and environmental costs because it directly affects the fuel efficiency. Effective intervention that delays flow transition is thus critical because wall-bounded turbulent flows generate significantly higher surface friction drag compared to laminar flows. Flow separation, as a contrasting scenario, is key to overall aerodynamic/hydrodynamic stability, form drag, vehicle maneuverability, and applications that require turbulent mixing. Similar considerations for fluid-structure interaction are applicable to other problems, such as land vehicles, wind turbines, and long-range pipelines, to name a few.
@@
Laminar-to-turbulent transition in walled-bounded flows has been investigated extensively in the past, leading to a rich build-up of knowledge on how transition occurs across various flow regimes [3]. A “natural” form of transition is associated with the growth in amplitude of infinitesimal fluctuations (also referred to as perturbations or disturbances  [^1] These terms are used interchangeably in this article.) representing unstable propagating modes inherent in the flow. If left without intervention, these modes advance to nonlinear growth, secondary transition mechanisms, flow structure breakdown, and finally fully developed turbulent flow [3]. A dominant primary mode in air or water channel flows or boundary layers is known as Tollmien-Schlichting (T-S) waves [5, 6]. These unstable waves, which travel with the mean flow, have a vortical nature and grow over an identifiable narrow band of frequencies that can be predicted with linear stability analysis [7, 8]. Although T-S waves are not always the mechanism directly responsible for transition, particularly under realistic flight conditions, much research has been devoted to their understanding [9] due to their fundamental nature, and because they serve as a platform for development of new technologies for laminar flow control involving more complex types of instabilities.
@@
Figure 1: Illustration of scatterless interferences: (a) Classical scattering of acoustic waves by a lattice of elastic objects, where the interfering waves and the objects are in the same space. (b) Scatterless interferences of flow fluctuation waves by a lattice of PSubs, where the waves and the PSubs are not in the same space.
@@
Numerous research approaches have been proposed in the last couple of decades aimed at passively mitigating the undesirable effects of fluid-structure interactions. These include installation of riblets over the surface [10, 11], creation of artificial surface roughness [12, 13], drilling of holes to form a porous surface [14, 15], or coating the surface with a compliant (low stiffness) material [16, 17, 18, 19, 20, 21, 22]. None of these approaches, however, synchronize with the frequency, phase, and wavevector characteristics of the flow instabilities, and are therefore limited in their effectiveness. An ideal intervention requires a tailored solution, with mechanistic precision, to create a passive and responsive control stimulus that accounts for the dynamical properties of the underlying flow transition mechanisms. In 2015, the general concept of flow control by subsurface phonons has been introduced as an approach capable of achieving this level of precise wave-synchronized control of flow instabilities [23]. A phononic subsurface (PSub) comprises a synthetically designed architectured material affixed beneath  [^2] Installation of a PSub may be effectively concealed because it is placed underneath the surface. In other words, an external observer viewing a PSub-installed airplane wing would view a normally looking surface and would not be able to identify if a PSub is present or not. the surface exposed to the flow (e.g., of a wing or vehicle body). The function of the PSub is to manipulate small-amplitude vibrations on the surface, and by extension the flow fluctuations near the wall that are responsible for transition. PSubs may be designed to passively respond to the flow instabilities in an out-of-phase manner, creating stabilization, or in an in-phase manner, creating destabilization--either function is realized by design. In past computational investigations, a PSub has been applied as a solitary unit [23, 25, 26, 27, 28] or as a contiguous layout of units distributed along the streamwise direction [29, 30].
@@
Despite the rapid progress of research on flow control by subsurface phonons, critical aspects remain to be addressed for the general concept to reach its potential. Two key limitations of previous demonstrations are that the PSub effect is not effective downstream to the PSub location, and that it is applicable only on unidirectional instability waves. The former must be resolved to enable transition delay, and the latter is significant for real-world flow control. While the downstream control objective was addressed with a “multiple-input-multiple-output” PSub configuration, offering a remarkable explicit display of transition delay [25], that approach is inherently limited to unidirectional instabilities because it dictates that a single PSub must interact with more than one flow point requiring a phased relation to be tuned specifically along a certain direction. Furthermore, the fixed distance between the input/output points limits the approach to a narrow band of fluctuation wavenumbers. Schmidt et al. [28] explored the concept from an alternative angle: rather than engage with subsurface phonons to inhibit the fluctuation production mechanisms (as done in all previous PSub studies), they designed their PSub based on pass-band motion to absorb and trap the fluctuation energy by resorting to actively controlled time-varying material properties. This approach enhances the energy transfer from the mean flow to the instabilities, rather than reduces it [26]. When more energy is drawn from the mean flow, this energy will accumulate in the subsurface structure requiring its dissipation as heat.
@@
In this work, we present the concept of a lattice of PSubs comprising a collection of individual PSub units laid out following a square or hexagonal lattice symmetry  [^3] Other lattice symmetries may also be considered in the future.. Similar to acoustic or elastic waves propagating around or through a lattice of scatterers (see Fig. 1a), we design our configuration on the basis of a rigorous Bloch wave analysis where the field variable is the flow fluctuations. Yet, in our system there are no scatterers. In contrast to classical scattering problems where the interfering waves and the scattering objects are in the same space, the flow instability waves interfere as a result of the local influence of each PSub on the fluctuation velocity components. Here we recall that the PSub is beneath the space of the flow (see Fig. 1b). After developing Bloch’s theorem for this unique problem, we demonstrate by direct numerical simulations of the coupled fluid-Psub lattice system sustained downstream reduction in the fluctuations kinetic energy, which indicates a delay in laminar-to-turbulent transition. Moreover, as dictated by the PSub lattice symmetry, this approach is immune to any changes in the direction of propagation of the instability waves and may be tuned to accommodate a wide range of wavenumbers along each direction--similar to the classical acoustic Bragg scattering problem which we also demonstrate to provide a direct analogy and comparison.
@@ Results
We form our lattice of Psubs using identical PSub units, all installed in a channel model. We select a channel flow problem for simplicity, but all the underlying concepts are readily applicable to boundary layer flows. The Psub design used in this investigation is shown in Fig. 2a. It consists of an elastic rod, composed of five unit cells, with each unit cell accomodating a local resonator (which in practice may be realized as a cantilevered pillar). The dispersion curves and the amplitude and phase frequency-response characteristics for this nominal PSub configuration is shown in Fig. 2b. The product of the amplitude and phase yields a performance metric which provides an a priori prediction of the beahvior of the PSub once passively engaged with the flow--this quantity is also plotted in Fig. 2b.
@@
Figure 3 shows the results of applying Bloch analysis over a unit cell comprising the flow interacting with the PSub in three configurations, a full-span PSub, and the square and hexagonal lattice arrangements [see schematics (b), (c) and (d) in Fig. 3]. The PSub interaction with the flow has been accounted for through a frequency-dependent complex-valued admittance, denoted by $Y$ in Fig. 3(a). The admittance is obtained, for a given frequency, from the solution of the PSub system of equations [25] and it is closely related to the performance metric. We use this function alongside transpiration boundary conditions to find the most unstable eigenvalue--and associated eigenmode--from the standard Orr-Sommerfeld equations modified to accommodate the periodicity imposed by the application of Bloch’s theorem in terms of the T-S wave’s wavenumber, which corresponds to the eigenvalue. The mathematical details of the formulation can be found in the Supplementary Material. As a result of this analysis, the dispersion curves corresponding to the unstable flow mode in Fig. 3(a) have been obtained. These results are complemented by the rate of production of perturbation kinetic energy (PKE)--denoted by $P_{\text{r}}$ in Fig. 3--obtained from the Bloch analysis. These have been computed from the eigenmodes associated with the unstable eigenvalues, normalized by the average PKE computed for the rigid case (refer to Supplementary Materials for details). Results for both weak destabilizing and stabilizing cases, at frequencies of 1539.5 Hz and 1555 Hz respectively, have been averaged over a flow domain close to the PSub wall, in particular along 10% of the channel’s half height. Contour plots showing the $P_{\text{r}}$ spanwise distribution over three unit cells (in the streamwise direction) are given in Fig. 3(b) to (d) for the full-span and for each of the lattice configurations. The average $P_{\text{r}}$ values over a spanwise period are shown in the last row of Fig. 3(b) to (d) to summarize the results. The different cases are compared with the results for the rigid case, to show the spatial stabilizing/destabilizing effects when the flow interacts with each PSub patch.
@@
The ultimate test for any flow control approach is through high-fidelity simulations, particularly direct numerical simulations (DNS) of the Navier-Stokes equations, as these resolve the high spatial and temporal frequencies of the flow fluctuations--which are central to the transition mechanism. We execute simulations in a three-dimensional (3D) computational domain and retain all nonlinear terms. In Fig. 4 we present DNS results for the two lattice configurations, and also show the results of an analogous acoustic scattering simulation for comparison. The coupled fluid-structure simulations were run at $Re=7500$ with a 1555 Hz unstable T-S wave input at the left end of the channel. See the Models and Methods sections for more details on the model and simulation parameters and numerical techniques used. The results in Figs. 4b and 4c show a sustained reduction in the PKE downstram of the PSub lattice, with a slightly stronger effect by the hexagonal arrangement compared to the square arrangement.
@@ Discussion
Results from the Bloch analysis show a correlation between the dispersion curves (in Fig. 3) and the PSub performance metric (in Fig. 2). It is observed that the intensity of the T-S wave’s instability--in this case given by the imaginary wavenumber--rapidly grows in the vicinity of the resonance frequency, which helps enhance the stabilizing/destabilizing effects. Furthermore, the rate of production of PKE within the PSub region consistently increases in the destabilizing case and decreases--even reaching negative values--in the stabilizing case, in agreement with the stand-alone performance metric predictions (see Fig. 3(b) to (d)). Interestingly, the $P_{\text{r}}$ also exhibits an opposite effect upon leaving the PSub, which is in accordance with what we observe in the fluid-structure interaction (FSI) simulations (see, for instance, how the PKE slightly increases in Figs. 4(b) and (c) each time the flow leaves a PSub unit behind despite being stabilizing cases). This analysis also shows how the full-span case exhibits a stronger response than both lattice configurations, which demonstrates that spanwise interactions can affect the flow instabilities downstream, as demonstrated by the full-scale FSI simulations results. While both square and hexagonal PSub lattices are effective in sustaining downstream stabilization, both the Bloch analysis and simulations indicaete that the hexagonal configuration is slightly superior--its PKE slope in Fig. 4c is $\theta=4.6^{\circ}$ compared to $\theta=4.4^{\circ}$ for the square lattice. Upon further design optimization of both the PSub unit and the lattice configuration, stronger downstream reductions in PKE are attainable. It is clearly observed that both flow fluctuations “scatterless” interferences are nearly identical to the interferences triggered by acoustic scattering.
@@
Figure 2: PSub design and its dispersion and vibration response characteristics: (a) Schematic of locally-resonant elastic metamaterial-based PSub utilized in this study. The length of the unit cell is $a_{\rm Psub}=1$ cm yielding a total PSub length of 5 cm. Each PSub is installed in the flow subsurface with its top face directly exposed to the flow. Flow instabilities, e.g. T-S waves, will continuously excite the PSub at the top edge and the PSub, in turn, will respond at the same point. Depending on the frequency, the PSub will respond according to its dynamical characteristics. Close to an anti-resonance or a resonance, the PSub motion will respectively impede or enhance the energy extraction from the mean flow into the fluctuation field [26]. This passive process will repeat and cause sustained control of incoming instability waves near the wall at the location of the PSub. (b) Four key plots that characterize the PSubs dynamics. From top to bottom, the following are shown: dispersion curves for PSub unit cell, steady-state vibration amplitude and phase response of the PSub top edge when excited at the same location, and performance metric obtained by multiplying the amplitude by the phase. All plots are obtained by analyzing a stand-alone finite-element model of the PSub.
@@
Figure 3: (a) Dispersion curves for selected unstable flow mode in the presence of a PSub lattice. The first and second rows correspond to the imaginary and real components of the wavenumber $\kappa$ for the rigid cases and the PSub cases in the full-span, square-lattice and hexagonal-lattice configurations [depicted by the (b), (c) and (d) unit cell schematics, respectively]. The third row shows the real and imaginary components of the PSub’s admittance $Y$ at different frequencies. The dashed vertical lines indicate the frequencies for which the average modal perturbation kinetic energy (PKE) production rate, $P_{\text{r}}$, distributions on the right plots have been obtained. Results for each frequency are depicted in their corresponding panel. The first panel (in orange) shows a weak destabilizing case (at a frequency of 1539.5 Hz) and the second panel (in light green) is for a weak stabilizing case (at a frequency of 1555 Hz). The curves below the panels correspond to the averaged production rate evolution obtained for both the weak destabilizing and stabilizing cases compared to that of the rigid case. As a reference, the average has been evaluated in the vicinity of the PSub wall (in the range of 10% of half of the channel’s height), and all the modal amplitudes have been normalized to make them coincide with that of the rigid case at the left edge of the first unit cell.
@@
Figure 4: Demonstration of scatterless interferences by DNS: (a) Acoustic wave propagation in a space covered by a lattice of elastic scatterers, considering square symmetry (left) and hexagonal symmetry (right). (b) Time-averaged flow PKE as a function of the streamwise and spanwise directions for flow in a channel with PSubs laid out following both symmetries. (c) Time- and spanwise-averaged PKE as a function of the streamwise position. The results in (b) and (c) indicate interference patterns, even though the surface is flat and smooth and there are no objects in the flow space. These interferences stem for the collective action of the local effect by each PSub. Figure 4(c) shows time-averaged PKE on and off the centerline of the channel, reflecting the effects of the different symmetries of the square and hexagonal lattices. The time- and spanwise-averaged PKE is shown in dark green, where the angle $\theta$ at the end of the channel indicates a decrease in the slope of PKE compared to the rigid case, and hence transition delay downstream of the PSub lattice. The angle is measured as $\theta=4.4^{\circ}$ and $\theta=4.6^{\circ}$ for the square and hexagonal cases, respectively, indicating the transition is delayed further by the hexagonal PSubs lattice.
@@
In conclusion, we have demonstrated that by engineering the phonon properties of the subsurface, which could very well be concealed, it is possible to almost completely execute full command of the very nature of the fluid-structure interaction. Such intervention, which is fully passive, offers precise mechanistic control of the behavior of wall-bounded flows. Our findings demonstrate that the principles of Bloch analysis, traditionally applied to acoustic and elastic waves, can be effectively translated into the realm of flow control through the implementation of PSubs. By leveraging locally resonant metamaterial-based phononic subsurfaces, we have shown that it is possible to manipulate flow instabilities in a manner analogous to wave propagation in periodic acoustic and elastic media. This novel approach enables passive control of phenomena such as transition to turbulence, with lattice arrangements of PSubs facilitating attenuation of T-S waves’ instabilities downstream, while also being robust to variations in wavenumber and direction of propagation. With the advent of concepts from phonon engineering and metamaterials, the prospects of future improvements of the performance is highly accessible. It is conceivable, for example, that with architectured materials, a fully laminar boundary layer around an airplane may be attained in the near future. Such outcome would drastically improve fuel efficiency. Potentially, the Psub lattices may be designed to simultaneously inhibit noise emission as well, which would bring about the added advantage of passenger comfort and reduced noise pollution.

### v2

@@ Abstract
Turbulent boundary-layer flows generate significantly higher surface-friction drag than flows that remain laminar. Avoidance or delay of laminar-to-turbulent transition is therefore critical for the efficient operation of air and water vehicles. However, control of the flow perturbations that trigger this transition is difficult due to their unstable and persistent nature. Here we present a passive, phonons-based approach that targets these perturbations through engineered interactions at the flow-surface interface. This intervention is achieved via a tailored frequency- and wavevector-dependent coupling between the flow and a smooth, homogeneous surface, beneath which a functional material architecture is entirely embedded. The subsurface consists of a wall-parallel lattice of wall-normal phononic structural units. Each discrete phononic subsurface (PSub) unit is designed to respond to local flow disturbances in an out-of-phase manner, effectively suppressing the perturbations at the point of interaction. Collectively, the lattice induces interference effects that cause the kinetic energy of the flow instabilities to decay downstream of the intervention region, and thereby delaying the transition. To guide the design of the subsurface lattice, a Bloch-wave analysis is developed for the flow perturbations within a representative unit cell interfaced with a PSub unit. This is followed by direct numerical simulations to validate the concept. This approach represents a paradigm shift in the design philosophy for aircraft, ship, and submarine surfaces--moving beyond streamlined shaping to leveraging subsurface phonon engineering for the reduction of drag and enhancement of overall performance.
Significance
Airflow over aircraft wings or around the fuselage generates skin-friction drag, severely reducing fuel efficiency. Similar drag effects occur in water flow around ships and submarines and in the transport of gas or oil through long-range pipelines. This friction drag can be significantly reduced by maintaining laminar flow or delaying the transition to turbulence. In this work, we demonstrate that phonon motion—low-amplitude intrinsic vibrations—within a lattice of architected materials embedded just beneath the flow-exposed surface can be engineered to fundamentally alter the fluid-structure interaction. This, in turn, enables the suppression of the flow instabilities responsible for triggering the transition to turbulence. The phonon properties of each subsurface unit in the lattice are precisely tailored to influence the flow locally, while the collective arrangement is engineered to produce a favorable downstream interference effect within the flow. This approach may alternatively be adapted for flow destabilization to delay separation or to enhance chemical mixing and combustion, offering a versatile platform for a broad range of flow control applications.
Introduction
@@
The interaction between a flowing fluid and a surface is a dynamical process central to a broad range of both natural and artificial systems. Two most consequential examples are flow over an aircraft wing and around a ship’s hull. The nature of the fluid-structure interaction in these cases governs critical phenomena such as the intensity of skin-friction drag, the onset location of laminar-to-turbulent transition, and the likelihood of flow separation—all of which have significant implications for overall vehicle performance [1, 2]. Skin-friction drag, in particular, carries substantial economic and environmental costs, as it directly impacts fuel efficiency. Effective intervention that delays flow transition is thus vital because wall-bounded turbulent flows may generate an order-of-magnitude or more larger surface-friction drag compared to laminar flows. On the other hand, flow separation influences aerodynamic and hydrodynamic stability, form drag, and vehicle maneuverability. Targeted intervention is therefore highly beneficial for control of flow separation in these applications. Similar fluid-structure-interaction challenges are encountered in various other contexts, including land vehicles, wind turbines, long-range pipelines, processes requiring enhanced turbulent mixing, among others.
@@
The laminar-to-turbulent transition in wall-bounded flows has been the subject of extensive research, resulting in a rich build-up of knowledge on how transition unfolds across different flow regimes [3]. A “natural” transition pathway typically involves the growth of infinitesimal flow perturbations--also referred to as disturbances  [^1] These terms are used interchangeably in the literature.--which represent inherent unstable modes within the flow. Without intervention, these modes may advance to nonlinear amplification, leading to flow structure breakdown and the ultimate evolution into fully developed turbulence [3]. A dominant primary mode in air or water channel flows, as well as boundary layers, is known as the Tollmien-Schlichting (T-S) wave [5, 6]. Unstable T-S waves take a vorticial form and travel with the mean flow, growing over an identifiable narrow band of frequencies that can be predicted with linear stability analysis [7, 8]. While T-S waves are not always the direct cause of transition, especially under complex or realistic operating conditions, they have been the focus of extensive research [9]. This is due to their fundamental nature, and because they serve as a platform for development of new technologies for laminar flow control involving more complex types of instabilities.
@@
Figure 1: Illustration of scatterless interferences: (a) Classical scattering of acoustic waves by a lattice of elastic objects, where the interfering waves and the objects are in the same space. (b) Scatterless interferences of flow perturbations waves by a lattice of PSubs, where the waves and the PSubs are not in the same space.
@@
Numerous research approaches have been proposed in the last couple of decades aimed at passively mitigating the undesirable effects of fluid-structure interactions. These include installation of riblets over the surface [10, 11], creation of artificial surface roughness [12, 13], drilling of holes to form a porous surface [14, 15], or coating the surface with a compliant (low-stiffness) material [16, 17, 18, 19, 20, 21, 22]. These approaches, however, are not based on a synchronization with the frequency, phase, and wavevector characteristics of the flow instabilities, and are therefore limited in their effectiveness. An ideal intervention requires a tailored solution, with mechanistic precision, to create a passive and responsive control stimulus that accounts for the dynamical properties of the underlying flow transition mechanisms. In 2015, the general concept of flow control by subsurface phonons was introduced as an approach capable of achieving this level of precise wave-synchronized control of flow instabilities [23]. A phononic subsurface (PSub) comprises a synthetically designed architected material affixed beneath  [^2] Installation of a PSub may effectively be concealed because it is placed underneath the surface. In other words, an external observer viewing a PSub-installed airplane wing would view a normally looking surface and would not be able to identify if a PSub is present or not. the surface exposed to the flow (e.g., of a wing or vehicle body). The function of the PSub is to manipulate small-amplitude vibrations on the surface, and by extension the flow perturbations near the wall that are responsible for transition. PSubs may be designed to passively respond to the flow instabilities in an out-of-phase manner, creating stabilization, or in an in-phase manner, creating destabilization--either function is realized a priori by design, or in the future may be switched or tuned in real time by actuation. In past computational investigations, a PSub has been applied as a solitary unit [23, 25, 26, 27, 28] or as a contiguous layout of units distributed along the streamwise direction [29, 30].
@@
Despite the rapid progress of research on flow control by subsurface phonons, critical aspects remain to be addressed for the general concept to reach its potential. Two key limitations of previous demonstrations are that the PSub effect is not effective downstream to the PSub location, and that it is applicable only to unidirectional instability waves. The former must be resolved to enable transition delay, and the latter is significant for real-world flow control--where cross-flow scenarios may be encountered, for example. While the downstream control objective was addressed with a “multiple-input-multiple-output” PSub configuration, offering a remarkable explicit display of transition delay [25], that approach is inherently limited to unidirectional instabilities because it dictates that a single PSub must interact with more than one flow point requiring a phased relation to be tuned specifically along a certain direction. Furthermore, the fixed distance between the input/output points limits the approach to a narrow band of perturbation wavenumbers. Schmidt et al. [28] explored the concept from an alternative angle: rather than engage with subsurface phonons to inhibit the perturbation production mechanisms (as done in all previous PSub studies), they designed their PSub based on pass-band motion to absorb and trap, by filtering, the energy of undesirable fluctuations. Actively controlled time-varying material properties were used for demonstration, however the principle may be implemented passively using nonlinear effects. This approach, which is applicable to only the function of stabilization, permanently attracts the energy that is continuously being transferred from the mean flow to the instabilities, instead of reducing it at the source [26]. This energy will accumulate in the subsurface structure requiring its ultimate dissipation as heat.
@@
In this work, we present the concept of a lattice of PSubs comprising a collection of individual PSub units laid out following a square or hexagonal lattice symmetry  [^3] Other lattice symmetries may also be considered in the future.. Similar to acoustic or elastic waves propagating around or through a lattice of rigid or elastic scatterers, respectively (see Fig. 1a), we design our configuration on the basis of a rigorous Bloch wave analysis where the field variable is the flow perturbations. Yet, in our system there are no scatterers. In contrast to classical scattering problems where the interfering waves and the scattering objects are in the same space, the flow instability waves interfere as a result of the local influence of each PSub on the perturbation velocity components. Here we recall that the PSubs are located beneath the spatial domain of the flow (see Fig. 1b). After developing Bloch’s theorem for this unique problem, we demonstrate by direct numerical simulation (DNS) of the coupled fluid-PSub lattice systems sustained downstream reduction in the perturbation kinetic energy (KE), which indicates a delay in laminar-to-turbulent transition. Moreover, as dictated by the PSub lattice symmetry, this approach is immune to any changes in the direction of propagation of the instability waves and may be tuned to accommodate a relatively wide range of wavenumbers along each direction--similar to the classical acoustic Bragg scattering problem which we also demonstrate to provide a direct analogy and comparison.
@@ Results
We form our lattice of PSubs using identical PSub units, all installed in a channel model. We select a channel flow problem for simplicity, but all the underlying concepts are readily applicable to boundary layer flows. The design of the PSub unit used in this investigation is shown in Fig. 2a. It consists of an elastic rod, composed of five unit cells, with each unit cell accommodating a local resonator (which in practice may be realized as a cantilevered pillar).The dispersion curves and the amplitude and phase frequency-response characteristics for this nominal PSub configuration is shown in Fig. 2b. The product of the amplitude and phase yields a performance metric [23] which provides an a priori prediction of the behavior of the PSub once passively engaged with the flow--this quantity is also plotted in Fig. 2b.
@@
Figure 3 shows the results of applying Bloch analysis over a unit cell comprising the flow interacting with the PSub in three configurations, a full-span PSub, and the square and hexagonal lattice arrangements [see schematics (b), (c) and (d) in Fig. 3]. The PSub interaction with the flow has been accounted for through a frequency-dependent complex-valued admittance, denoted by $Y$ in Fig. 3(a). The admittance is obtained, for a given frequency, from the solution of the PSub system of equations [25] and it is closely related to the performance metric. We use this function alongside transpiration boundary conditions to find the most unstable eigenvalue--and associated eigenmode--from the standard Orr-Sommerfeld equations modified to accommodate the periodicity imposed by the application of Bloch’s theorem in terms of the T-S wave’s wavenumber, which corresponds to the eigenvalue. The mathematical details of the formulation and definitions of key quantities can be found in the Supplementary Material. As a result of this analysis, the dispersion curves corresponding to the unstable flow mode in Fig. 3(a) have been obtained. These results are complemented by computing the average perturbation kinetic energy, denoted by $K_{\text{p}}$, and the rate of production of that energy, denoted by $P_{\text{r}}$. The perturbation KE has been computed from the eigenmodes associated with the unstable eigenvalues, normalized by the corresponding value for the rigid case. The $P_{\text{r}}$ quantity, as obtained from the Bloch analysis, is plotted in Fig. 3b-c. The reader is referred to Supplementary Materials for calculation details. Results for both a weak destabilizing case and a weak stabilizing case, at frequencies of 1539.5 Hz and 1555 Hz respectively, have been averaged over a flow domain close to the PSub wall, in particular along 10% of the channel’s half height. Contour plots showing the $P_{\text{r}}$ spanwise distribution over three unit cells (in the streamwise direction) are given in Fig. 3(b) to (d) for the full-span and for each of the lattice configurations. The average $P_{\text{r}}$ values over a spanwise period are shown in the last row of Fig. 3(b) to (d) to summarize the results. The different cases are compared with the results for the rigid case, to show the spatial stabilizing/destabilizing effects when the flow interacts with each PSub unit.
@@
The ultimate test for any flow control approach is through high-fidelity simulations, particularly DNS of the Navier-Stokes equations, as these resolve the high spatial and temporal frequencies of the flow perturbations--which are central to the transition mechanism. We execute simulations in a three-dimensional (3D) computational domain and retain all nonlinear terms. In Fig. 4 we present DNS results for the two lattice configurations, and also show the results of an analogous acoustic scattering simulation for comparison. The coupled fluid-structure simulations were run at $Re=7500$ with a 1555 Hz unstable T-S wave input at the left end of the channel. See the Models and Methods sections for more details on the model, simulation parameters, and numerical techniques used. The results in Figs. 4b and 4c show a sustained reduction in $K_{\text{p}}$ downstream of the PSub lattice, with a slightly stronger effect by the hexagonal arrangement compared to the square arrangement. This behavior is indicative of a delay in laminar-to-turbulent transition downstream of the PSub lattice region.
@@ Discussion
Results from the Bloch analysis show a correlation between the dispersion curves (in Fig. 3) and the PSub performance metric (in Fig. 2). It is observed that the intensity of the T-S wave’s instability--in this case given by the imaginary wavenumber--rapidly grows in the vicinity of the resonance frequency, which helps enhance the stabilizing/destabilizing effects. Furthermore, the rate of production $P_{\text{r}}$ within the PSub region consistently increases in the destabilizing case and decreases--even reaching negative values--in the stabilizing case, in agreement with the stand-alone performance metric predictions (see Fig. 3(b) to (d)). Interestingly, $P_{\text{r}}$ also exhibits an opposite effect at the trailing edge of each PSub unit, which is in accordance with what we observe in the fluid-structure interaction (FSI) simulations (see, for instance, how the $K_{\text{p}}$ slightly increases in Fig. 4(b), for both lattice configurations, each time the flow “leaves” a PSub unit behind despite being stabilizing cases). This analysis also shows that the leading and trailing edge changes in $P_{\text{r}}$ are stronger in the full-span case compared to the two lattice configurations, except that this case does not experience spanwise interactions. While both square and hexagonal PSub lattices are effective in sustaining downstream stabilization, both the Bloch analysis and simulations indicate that the hexagonal configuration is slightly superior--its $K_{\text{p}}$ slope in Fig. 4c is $\theta=4.6^{\circ}$ compared to $\theta=4.4^{\circ}$ for the square lattice. Upon further design optimization of both the PSub unit and the lattice configuration, stronger downstream reductions in $K_{\text{p}}$ are attainable. It is intriguingly observed that the flow perturbations “scatterless” interferences resemble the interferences triggered by acoustic scattering, for both lattice configurations.
@@
Figure 2: PSub design and its dispersion and vibration response characteristics: (a) Schematic of locally-resonant elastic metamaterial-based PSub unit utilized in this study. The length of the unit cell is $a_{\rm Psub}=1$ cm yielding a total PSub length of 5 cm. Each PSub is installed in the flow subsurface with its top face directly exposed to the flow. Flow instabilities, e.g. T-S waves, will continuously excite the PSub at the top edge and the PSub, in turn, will respond at the same point. Depending on the frequency, the PSub will respond according to its dynamical characteristics. Following an out-of-phase or in-phase response, the PSub motion will respectively impede or enhance the energy extraction from the mean flow into the perturbation field [26]. This passive process will repeat and cause sustained control of incoming instability waves near the wall at the location of the PSub. (b) Four key plots that characterize the PSubs dynamics. From top to bottom, the following are shown: dispersion curves for PSub unit cell, steady-state vibration amplitude and phase response of the PSub top edge when excited at the same location, and performance metric obtained by multiplying the amplitude by the phase. All plots are obtained by analyzing a stand-alone finite-element model of the PSub.
@@
Figure 3: (a) Dispersion curves for selected unstable flow mode in the presence of a PSub lattice. The first and second rows correspond to the imaginary and real components of the wavenumber $\kappa$ for the rigid-wall case and the PSub cases in the full-span, square-lattice, and hexagonal-lattice configurations [depicted by the (b), (c) and (d) unit cell schematics, respectively]. The third row shows the real and imaginary components of the PSub’s admittance $Y$ at different frequencies. The dashed vertical lines indicate the frequencies for which the average modal perturbation KE production rate, $P_{\text{r}}$, distributions on the right plots have been obtained. Results for each frequency are depicted in their corresponding panel. The first panel (in orange) shows a weak destabilizing case (at a frequency of 1539.5 Hz) and the second panel (in light green) is for a weak stabilizing case (at a frequency of 1555 Hz). The curves below the panels correspond to the averaged production rate evolution obtained for both the weak destabilizing and weak stabilizing cases compared to that of the rigid case. As a reference, the average has been evaluated in the vicinity of the PSub wall (in the range of 10% of half of the channel’s height), and all the modal amplitudes have been normalized to make them coincide with that of the rigid-wall case at the left edge of the first unit cell.
@@
Figure 4: Demonstration of scatterless interferences of flow perturbations by DNS: (a) Acoustic wave propagation in a space covered by a lattice of elastic scatterers, considering square symmetry (left) and hexagonal symmetry (right). (b) Time-averaged flow perturbation KE as a function of the streamwise and spanwise directions for flow in a channel with PSubs laid out following both symmetries. (c) Time- and spanwise-averaged perturbation KE as a function of the streamwise position. The results in (b) and (c) indicate interference patterns, even though the surface is flat and smooth and there are no objects in the flow space. These interferences stem for the collective action of the local effect by each PSub. Figure 4(c) shows time-averaged perturbation KE on and off the centerline of the channel, reflecting the effects of the different symmetries of the square and hexagonal lattices. The time- and spanwise-averaged perturbation KE is shown in dark green, where the angle $\theta$ at the end of the channel indicates a decrease in the slope of $K_{\text{p}}$ compared to the rigid-wall case, and hence transition delay downstream of the PSub lattice. The angle is measured as $\theta=4.4^{\circ}$ and $\theta=4.6^{\circ}$ for the square and hexagonal cases, respectively, indicating the transition is delayed further by the hexagonal PSubs lattice.
@@
In conclusion, we have shown that by engineering the phonon properties of the subsurface--virtually entirely concealed beneath the flow-exposed surface--it is possible to exert full command over the very nature of the fluid-structure interaction. This passive intervention offers precise mechanistic and tunable control of the behavior of wall-bounded flows. Our findings demonstrate that the principles of Bloch wave analysis, traditionally applied to electronic, photonic, acoustic, and elastic waves [32, 33], can be effectively extended to flow perturbations influenced by PSubs. By leveraging locally resonant metamaterial-based individual PSub units, we have shown that flow instabilities can be manipulated through interference effects, in a manner analogous to wave phenomena in periodic elastic or acoustic media. This approach enables passive control over critical processes such as the transition to turbulence, with the PSub lattice facilitating attenuation of unstable T-S waves downstream, while being robust to variations in wavenumber and direction of propagation. Other types of flow instabilities may be similarly treated in future studies. With the ongoing influx of new concepts from phonon engineering and metamaterials [34, 35, 36], the prospects of further future improvements of the performance following this framework is highly accessible, and so is the path to practical implementation given the rapid advancements in fabrication technology [37, 38]. A wide range of flow-control applications stand to benefit, spanning from hypersonic airflows to low-speed liquid transport. It is conceivable that within this framework future aircrafts will maintain a fully laminar boundary layer over their surfaces, enabling transformative improvements in fuel efficiency. In addition, PSub lattices may be simultaneously designed to inhibit aeroacoustic emissions, offering the additional rewards of improved passenger comfort and reduced environmental noise pollution. Similar advances will materialize for hydroacoustic emissions as well.